a certain rectangle is 3cm longer than it is wide. The area is 550. Find the dimensions

Answer by checkley77(12844)

W=L+3
LW=550
L(L+3)=550
L^2+3L=550
L^2+3L-550=0
(L+25)(L-22)=0
LL-22=0
L=22 ANS. FOR THE LENGTH.
W=22+3=25 ANS. FOR THE WIDTH.
PROOF:
22*25=550
550=550